# Nonstabilizerness of a Boundary Time Crystal

### v1

@@ Abstract
Boundary time crystals exhibit measurement-induced phase transitions in their steady-state entanglement, with critical behavior that depends on the particular unraveling of the Lindblad dynamics. In this work, we investigate another key measure of quantum complexity—nonstabilizerness (or “magic”)—and show that it follows a markedly different pattern. Importantly, in contrast to entanglement, for large system sizes, magic remains invariant under different unraveling schemes—a property we attribute to the inherent permutational symmetry of the model. Although the steady-state stabilizer entropy does not display a genuine phase transition, it exhibits a singular derivative (a cusp) at the mean-field critical point. Furthermore, we demonstrate that finite-size simulations of the average Lindblad evolution fail to capture the asymptotic behavior of magic in the time-crystal phase, while quantum trajectory unravelings correctly reveal its extensive scaling with system size. These findings offer insights into how different quantum resources manifest in open systems.
@@ I Introduction
Collective spin models describe ensembles of two-level systems, or qubits, in which all particles experience the same global dynamics due to collective interactions, leading to permutationally invariant quantum states. These models are particularly relevant in quantum optics, where atoms confined in a cavity interact through a shared electromagnetic field [1, 2, 3, 4, 5, 6, 7]. A prominent example is the boundary time crystal (BTC), a driven-dissipative many-body system that exhibits persistent oscillations, breaking time-translation symmetry [8]. BTCs arise from the interplay between coherent driving and collective dissipation and have been proposed as promising platforms for investigating nonequilibrium quantum phenomena, leading to a flourishing research field [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29]. In particular, they are known to host measurement-induced phase transitions (MIPTs) [30, 31, 32, 33, 34, 35, 36], where their entanglement exhibits a transition as a function of the dissipation rate, or measurement strength [37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57]. This naturally raises the question of whether other quantum complexity measures, beyond entanglement, might also undergo phase transitions in monitored BTCs.
@@
While the role of entanglement in characterizing quantum phases and dynamics is well established [80, 81, 82, 52], nonstabilizerness in many-body physics remains less explored [83], especially in the context of phase transitions, where only a handful of results are known [84, 85, 86, 87, 88, 89, 90]. Investigating nonstabilizerness in monitored many-body systems, such as BTCs, can provide insights into quantum complexity in open dynamics and potentially reveal new universal features of measurement-induced transitions. In particular, understanding whether magic exhibits a transition similar to that of entanglement in monitored BTCs could shed light on how different forms of quantum complexity behave under nonunitary evolution.
@@
Motivated by this, our goal is to study the monitored dynamics of a model admitting a boundary time crystal phase, i. e., a permutationally invariant system composed of $N$ qubits placed in an optical cavity, driven by an external laser field, and subject to collective Markovian decay [8, 6, 22, 23]. The unconditional dynamics of the system (i. e., the average over many independent experiments) is described by a master equation in the Lindblad form, which is able to capture a superradiant transition in the steady state. However, when coupled to a suitable measuring device that monitors and keeps track of environmental transitions induced by the jump operator (i. e., the evolution conditioned to the measurement outcomes), the dynamics can be described in terms of stochastic quantum trajectories [91, 92] that, on average, provide the Lindblad equation. Depending on the experimental apparatus, different measurement schemes—ranging from photodetection to homo/heterodyne detection—give rise to distinct unravelings of the same Lindbladian. While the evolution of any observable remains identical across different unravelings, nonlinear quantities such as entanglement and magic may depend on the specific measurement protocol [93, 94, 95, 52, 96].
@@
This superradiant model presents several advantages for investigating the monitored dynamics of magic. First, the reconstruction of quantum trajectories from experiments is feasible, as individual trajectories saturate within a logarithmic timescale in system size, thereby reducing postselection requirements [52]. Second, the model exhibits a measurement-induced phase transition in entanglement, which coincides with the superradiant phase transition in the average state [8, 6, 22, 52]. This naturally raises the question of whether magic undergoes a similar transition at the same critical point. Third, the symmetries of the model enable the application of efficient computational techniques for quantifying nonstabilizerness.
@@
Indeed, most existing studies on nonstabilizerness in many-body systems are constrained to small system sizes [84, 77, 86, 85, 97] or to systems that admit efficient classical representations, such as tensor networks or Gaussian states [98, 99, 87, 100, 101, 102, 103]. Among these, permutationally invariant states—such as those describing boundary time crystal states—constitute a particularly advantageous platform. Their high degree of symmetry allows for the efficient computation of magic measures, particularly those based on the Pauli spectrum, with polynomial rather than exponential scaling [104]. This renders them especially valuable for exploring nonstabilizerness in large systems [105]. Finally, despite its growing relevance in the Noisy Intermediate-Scale Quantum (NISQ) era, dominated by noise and environmental interactions, the role of magic in open and monitored quantum systems remains largely unexplored [87, 105, 90, 97].
@@
We first find that evaluating magic across trajectories yields an average that, in the thermodynamic limit, becomes independent of the specific unraveling protocol. This result can be shown analytically using a replica approach and relies on the fact that trajectory correlations become negligible in the thermodynamic limit when computing expectation values. This contrasts sharply with another measure of quantum complexity—the entanglement entropy—which exhibits a significant dependence on the unraveling scheme when averaged over trajectories [52].
@@
Approaching the thermodynamic limit, we observe that the time average of the trajectory-averaged magic is well captured by mean-field (MF) theory, provided the mean-field results are averaged over time and many random initial conditions. The stochasticity introduced by individual trajectories is effectively described by the average over random initial states in the MF limit. However, this agreement holds only for time averages, not for individual time traces, and manifests as a cusp at the transition to the time-crystal phase. Furthermore, the magic scales linearly with the system size $N$ asymptotically, consistent with previous findings in other monitored systems [97]. This indicates the absence of magic phase transitions in the model, although remnants of critical behavior are still visible in the cusp.
@@
Interestingly, evaluating magic on the Lindblad-averaged density matrix yields a time average that is not well described by the mean-field theory in the boundary time crystal phase for large system sizes. In this phase, the average density matrix remains mixed and does not approach a pure state in the thermodynamic limit, leading to the failure of mean-field theory in describing magic. This highlights that, while correlations are negligible when computing expectation values of observables, they still affect the mixed state in the large-size limit, as revealed by the behavior of magic.
@@
The rest of this manuscript is organized as follows. In Sec. II, we describe the superradiant model and the monitoring protocols adopted in our analysis. In Sec. III, we recall the definition of the magic measure considered in this paper, the stabilizer 2-Rényi entropy, and briefly discuss how to compute it efficiently in the case of permutationally invariant quantum states. We show also that the stabilizer entropy is mostly unraveling-independent for large system sizes. In Sec. IV, we discuss the thermodynamic limit of the model, which is exactly captured by mean-field theory. In Sec. V, we turn our attention to finite-size systems and study the dynamics of magic of the average state and compare the magic averaged over trajectories and time with the mean-field results averaged over time and initial conditions finding a good agreement. There is no agreement with the Lindblad average density matrix, which remains mixed also in the thermodynamic limit, a fact related to the presence of correlations. In Sec. VI, we draw our conclusions.
@@ II.1 Quantum Jumps
Thus, the update rule to first order in $\delta t$ is
@@ II.2 Quantum State Diffusion
An alternative approach is provided by quantum state diffusion (QSD), which is associated with a heterodyne detection scheme wherein the quadratures of the bosonic field are continuously measured [93]. In the QSD formalism, each pure state evolves according to the stochastic differential equation
@@ II.3 General Unravelings
The QJ and QSD schemes are two limiting cases of a more general detection protocol in which the system is coupled to a local oscillator with frequency $\mathrm{Re}(\mu)$, where $\mu\in\mathbb{C}$  [^1] The imaginary part $\mathrm{Im}(\mu)$ is related to phase-sensitive attenuation or amplification in the detection process.. In this more general framework, the jump operator and the Hamiltonian are shifted as follows:
@@
The flexibility in the choice of unraveling protocol enables a detailed investigation into how measurement schemes and the associated stochastic dynamics affect the quantum properties of the system. While linear observables remain invariant across different unravelings, nonlinear measures such as entanglement and magic can exhibit different behaviors under different unraveling protocols, providing deeper insights into the role of monitoring in open quantum systems.
@@ III Stabilizer entropy
Stabilizer entropies are commonly used to quantify quantum magic and are magic monotones for Rényi indices $k\geq 2$ [110]. For mixed states, the 2-Rényi stabilizer entropy (SRE) is expressed as
@@
where $\hat{P}$ are all the Pauli strings of the Pauli group $\mathcal{P}_{N}$ of $N$ qubits. One can also define the density of magic $m_{2}=\mathcal{M}_{2}/N$. For pure states $\hat{\rho}=\ket{\psi}\bra{\psi}$ |, |$\ket{\psi}$ is a stabilizer state if and only if $\hat{\rho}$ has precisely $2^{N}$ nonzero components in the operator basis given by the Pauli strings, while all other components are zero [72]. As such, the SRE provides a measure of the entropy associated with the probability distribution $\Pi_{P}=\mathrm{Tr}(\hat{P}\hat{\rho})^{2}/2^{N}$, adjusted by the entropy of a stabilizer state. If a mixed state can be produced by partially tracing a stabilizer pure state, it has zero magic. The SRE obeys the bound $0\leq\mathcal{M}_{2}<N\ln 2$, where $\mathcal{M}_{2}=0$ if and only if the state is a stabilizer.
@@ III.2 Nonstabilizerness and permutational invariance
Figure 1: (a) Magic density, (b) entanglement and (c) magnetization dynamics for different unravelings ($N=40$, $\omega_{0}=2\kappa$, averaged over $M=2000$ trajectories). The dotted line marks $\kappa t=8$: the distributions of magic, entanglement and magnetization of the three unravelings at this time are shown in Appendix A.
@@
Other measures of quantum complexity, however, might be more sensitive than $\mathcal{M}_{2}$ to residual correlations at finite size, leading to a stronger dependence on the monitoring protocol. A notable example is the bipartite entanglement entropy, which, for a pure state, is an entanglement monotone defined as
@@
Ref. [52] analyzed the entanglement properties of the superradiant model in Eq. (2) across its phase diagram, showing that the entanglement entropy shows significant differences between unravelings, particularly in its scaling behavior in the low-dissipative regime. In the QJ unraveling, it transitions from an area law to a $\log(\log N$) scaling, with a $\log N$ peak at a critical point. In contrast, in the QSD unraveling, it undergoes an area-to-area transition with a logarithmic peak at the same critical point. This stark difference arises because entanglement entropy is a nonlinear function of the quantum state, making it intrinsically sensitive to correlations, whereas stabilizer entropy, being expressible solely in terms of expectation values, is not.
To illustrate this point, we show one example of dynamics in Fig. 1 for a system of $N=40$ spins, in the BTC phase [$\omega_{0}=2\kappa$ in Eq. (2)]. The Lindblad dynamics is unraveled in terms of pure-state quantum trajectories, and averages are computed over $M=2000$ realizations for all considered unravelings, specified by the value of the parameter $\mu$ in Eq. (5). In panel (a), where we plot the magic density, the curves relative to the different unravelings are very close to each other (we have numerically verified that the difference between curves decreases by increasing $N$), whereas in panel (b) we see that the unravelings have completely different average bipartite entanglement. As expected, the magnetization along the $z$ axis, defined as $m_{z}=\langle\hat{S}_{z}\rangle/S$, is the same among all unravelings, panel (c), as for any other observable. Similar results are valid for the full distributions along quantum trajectories, as we show in Appendix A.
These findings highlight that different measures of quantum complexity capture distinct physical aspects of monitored systems. While stabilizer entropy remains largely unraveling-independent due to its reliance on expectation values, entanglement entropy—being inherently sensitive to correlations—exhibits a strong dependence on the monitoring protocol. This contrast highlights the fact that quantum complexity manifests in fundamentally different ways depending on the chosen metric, as each assigns a different weight to correlations in the system.

### v3

@@ Abstract
Boundary time crystals exhibit measurement-induced phase transitions in their steady-state entanglement, with critical behavior that depends on the particular unraveling of the Lindblad dynamics. In this work, we investigate another key measure of quantum complexity—nonstabilizerness (or “magic”)—and show that it follows a markedly different pattern. Importantly, in contrast to entanglement, for large system sizes, nonstabilizerness remains invariant under different unraveling schemes—a property we attribute to the inherent permutational symmetry of the model. Although the steady-state stabilizer entropy does not display a genuine phase transition, it exhibits a singular derivative (a cusp) at the mean-field critical point. Furthermore, we demonstrate that finite-size simulations of the average Lindblad evolution fail to capture the asymptotic behavior of nonstabilizerness in the time-crystal phase, while quantum trajectory unravelings correctly reveal its extensive scaling with system size. These findings offer insights into how different quantum resources manifest in open systems.
@@ I Introduction
Collective spin models describe ensembles of two-level systems, or qubits, in which all particles experience the same global dynamics due to collective interactions, leading to permutationally invariant quantum states. These models are particularly relevant in quantum optics, where atoms confined in a cavity interact through a shared electromagnetic field [1, 2, 3, 4, 5, 6, 7]. A prominent example is the boundary time crystal (BTC), a driven-dissipative many-body system that exhibits persistent oscillations, breaking continuous time-translation symmetry [8]. BTCs arise from the interplay between coherent driving and collective dissipation and have been proposed as promising platforms for investigating nonequilibrium quantum phenomena, leading to a flourishing research field [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29]. In particular, they are known to host measurement-induced phase transitions (MIPTs) [30, 31, 32, 33, 34, 35, 36], where their entanglement exhibits a transition as a function of the dissipation rate, or measurement strength [37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57]. This naturally raises the question of whether other quantum complexity measures, beyond entanglement, might also undergo phase transitions in monitored BTCs.
@@
While the role of entanglement in characterizing quantum phases and dynamics is well established [80, 81, 82, 52], nonstabilizerness in many-body physics remains less explored [83], especially in the context of phase transitions, where only a handful of results are known [84, 85, 86, 87, 88, 89, 90]. Investigating nonstabilizerness in monitored many-body systems, such as BTCs, can provide insights into quantum complexity in open dynamics and potentially reveal new universal features of measurement-induced transitions. In particular, understanding whether nonstabilizerness exhibits a transition similar to that of entanglement in monitored BTCs could shed light on how different forms of quantum complexity behave under nonunitary evolution.
@@
Motivated by this, our goal is to study the monitored dynamics of a model admitting a boundary time crystal phase, i. e., a permutationally invariant system composed of $N$ qubits placed in an optical cavity, driven by an external laser field, and subject to collective Markovian (memoryless) decay [8, 6, 22, 23]. The unconditional dynamics of the system (i. e., the average over many independent experiments) is described by a master equation in the Lindblad form, which is able to capture a superradiant transition in the steady state, where transition processes are described by the action of a “jump operator” acting on the system state. When coupled to a suitable device that monitors and keeps track of environmental transitions induced by the measurements (i. e., the evolution conditioned to the measurement outcomes), the dynamics can be described in terms of stochastic quantum trajectories [91, 92] that, only on average, provide the Lindblad equation. Depending on the experimental apparatus, different measurement schemes—ranging from photodetection, where one counts the number of photons escaping the cavity, to homo/heterodyne detection, where one instead measures the quadratures of the outgoing light field by interfering it with a reference signal—give rise to distinct unravelings of the same Lindbladian. While the evolution of any observable remains identical across different unravelings, nonlinear quantities such as entanglement and nonstabilizerness may depend on the specific measurement protocol [93, 94, 95, 52, 96].
@@
This superradiant model presents several advantages for investigating the monitored dynamics of nonstabilizerness. First, the reconstruction of quantum trajectories from experiments is feasible, as individual trajectories saturate within a logarithmic timescale in system size, thereby reducing postselection requirements [52]. Second, the model exhibits a measurement-induced phase transition in entanglement, which coincides with the superradiant phase transition in the average state [8, 6, 22, 52]. This naturally raises the question of whether nonstabilizerness undergoes a similar transition at the same critical point. Third, the symmetries of the model enable the application of efficient computational techniques for quantifying nonstabilizerness.
@@
Indeed, most existing studies on nonstabilizerness in many-body systems are constrained to small system sizes [84, 77, 86, 85, 97] or to systems that admit efficient classical representations, such as sign-problem-free Hamiltonians [98], tensor networks or Gaussian states [99, 100, 101, 87, 102, 103, 104, 105, 106, 107]. Among these, permutationally invariant states—such as those describing boundary time crystal states—constitute a particularly advantageous platform. Their high degree of symmetry allows for the efficient computation of magic measures, particularly those based on the Pauli spectrum (i. e., the expectation values of all multi-qubit Pauli operators), with polynomial rather than exponential scaling [108]. This renders them especially valuable for exploring nonstabilizerness in large systems [109]. Finally, despite its growing relevance in the Noisy Intermediate-Scale Quantum (NISQ) era, dominated by noise and environmental interactions, the role of nonstabilizerness in open and monitored quantum systems remains largely unexplored [87, 109, 90, 97].
We emphasize that although the nonstabilizerness is a quantum complexity measure important for quantum computation, and existing quantum computers have only few qubits, it is important to study it in a many-body context as we are doing here. In some sense the situation is similar to the one of the entanglement entropy. Also this quantity was introduced as a measure of quantum complexity for quantum computation, and then its scaling with the system size has been found to have a relevant role in quantum many-body physics, in the context of ground-state quantum phase transitions [80, 110], in distinguishing quantum thermalization from quantum integrable behavior [111, 112, 113, 114, 115, 116, 117, 118, 119, 120] and in the measurement-induced phase transitions [37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57]. In all these cases the scaling of the entanglement entropy with the system size was relevant to highlight behaviors appearing in the large-size limit. For the nonstabilizerness the situation is still at the beginning, but appears to be similar. For instance this quantity has already been used to characterize ground-state quantum phase transitions [121, 122, 123]. Our work provides a further contribution in this direction, focusing on a quantum phase transition involving the full monitored dynamics.
@@
We first find that evaluating nonstabilizerness across trajectories yields an average that, in the thermodynamic limit, becomes independent of the specific unraveling protocol. This result can be shown analytically using a replica approach and relies on the fact that trajectory correlations become negligible in the thermodynamic limit when computing expectation values. This contrasts sharply with another measure of quantum complexity—the entanglement entropy—which exhibits a significant dependence on the unraveling scheme when averaged over trajectories [52].
@@
Approaching the thermodynamic limit, we observe that the time average of the trajectory-averaged nonstabilizerness is well captured by mean-field (MF) theory, provided the mean-field results are averaged over time and many random initial conditions. The stochasticity introduced by individual trajectories is effectively described by the average over random initial states in the MF limit. However, this agreement holds only for time averages, not for individual time traces, and manifests as a cusp at the transition to the time-crystal phase. Furthermore, nonstabilizerness scales linearly with the system size $N$ asymptotically, consistent with previous findings in other monitored systems [97]. This indicates the absence of nonstabilizerness phase transitions in the model, although signs of critical behavior are still visible in the cusp.
@@
Interestingly, evaluating nonstabilizerness on the Lindblad-averaged density matrix yields a time average that is not well described by the mean-field theory in the boundary time crystal phase for large system sizes. In this phase, the average density matrix remains mixed and does not approach a pure state in the thermodynamic limit, leading to the failure of mean-field theory in describing nonstabilizerness. This highlights that, while correlations are negligible when computing expectation values of observables, they still affect the mixed state in the large-size limit, as revealed by the behavior of nonstabilizerness.
@@
The rest of this manuscript is organized as follows. In Sec. II, we describe the superradiant model and the monitoring protocols adopted in our analysis. In Sec. III, we recall the definition of the magic measure considered in this paper, the stabilizer 2-Rényi entropy, and briefly discuss how to compute it efficiently in the case of permutationally invariant quantum states. We show also that the stabilizer entropy is mostly unraveling-independent for large system sizes. In Sec. IV, we discuss the thermodynamic limit of the model, which is exactly captured by mean-field theory. In Sec. V, we turn our attention to finite-size systems and study the dynamics of the stabilizer entropy of the average state and compare the nonstabilizerness averaged over trajectories and time with the mean-field results averaged over time and initial conditions finding a good agreement. There is no agreement with the Lindblad average density matrix, which remains mixed also in the thermodynamic limit, a fact related to the presence of correlations. In Sec. VI, we draw our conclusions.
@@ II.1 Quantum Jumps
Thus, the update rule to first order in $\delta t$ is [93]
@@ II.2 Quantum State Diffusion
An alternative approach is provided by quantum state diffusion (QSD), which is associated with a heterodyne detection scheme wherein the quadratures of the bosonic field are continuously measured. In the QSD formalism, each pure state evolves according to the stochastic differential equation [93]
@@ II.3 General Unravelings
The QJ and QSD schemes are two limiting cases of a more general detection protocol in which the system is coupled to a local oscillator with frequency $\mathrm{Re}(\mu)$, where $\mu\in\mathbb{C}$  [^1] The imaginary part $\mathrm{Im}(\mu)$ is related to phase-sensitive attenuation or amplification in the detection process.. In this more general framework, the jump operator and the Hamiltonian are shifted as follows [93]:
@@
The flexibility in the choice of unraveling protocol enables a detailed investigation into how measurement schemes and the associated stochastic dynamics affect the quantum properties of the system. While linear observables remain invariant across different unravelings, nonlinear measures such as entanglement and nonstabilizerness can exhibit different behaviors under different unraveling protocols, providing deeper insights into the role of monitoring in open quantum systems.
@@ III Stabilizer entropy
Stabilizer entropies are commonly used to quantify nonstabilizerness and are magic monotones for Rényi indices $k\geq 2$ [99, 128]. For mixed states, the 2-Rényi stabilizer entropy (SRE) is expressed as [129]
@@
where $\hat{P}=\sigma_{1}^{\alpha_{1}}\otimes\dots\otimes\sigma_{N}^{\alpha_{N}}$ are all the Pauli strings that constitute the Pauli group $\mathcal{P}_{N}$ of $N$ qubits. One can also define the density of SRE $m_{2}=\mathcal{M}_{2}/N$. For pure states $\hat{\rho}=\ket{\psi}\bra{\psi}$ |, |$\ket{\psi}$ is a stabilizer state if and only if $\hat{\rho}$ has precisely $2^{N}$ nonzero components in the operator basis given by the Pauli strings, while all other components are zero [72]. As such, the SRE provides a measure of the entropy associated with the probability distribution $\Pi_{P}=\mathrm{Tr}(\hat{P}\hat{\rho})^{2}/2^{N}$, adjusted by the entropy of a stabilizer state. If a mixed state can be produced by partially tracing a stabilizer pure state, its SRE is zero. The SRE obeys the bound $0\leq\mathcal{M}_{2}<N\ln 2$, where $\mathcal{M}_{2}=0$ if and only if the state is a stabilizer.
@@ III.1 Unravelling-independence of stabilizer entropy with negligible correlations
Figure 1: (a) Nonstabilizerness density, (b) entanglement and (c) magnetization dynamics for different unravelings ($N=40$, $\omega_{0}=2\kappa$, averaged over $M=2000$ trajectories). The dotted line marks $\kappa t=8$: the distributions of nonstabilizerness, entanglement and magnetization of the three unravelings at this time are shown in Appendix A.
@@ III.2 Nonstabilizerness and permutational invariance
Other measures of quantum complexity, however, might be more sensitive than $\mathcal{M}_{2}$ to residual correlations at finite size, leading to a stronger dependence on the monitoring protocol. A notable example is the bipartite entanglement entropy, which, for a pure state, is an entanglement monotone defined as [75]
@@
Ref. [52] analyzed the entanglement properties of the superradiant model in Eq. (2) across its phase diagram, showing that the entanglement entropy shows significant differences between unravelings, particularly in its scaling behavior in the low-dissipative regime. In the QJ unraveling, it transitions from an area law, where the enetanglement entropy is independent of $N$, to a $\log(\log N$) scaling, with a $\log N$ peak at a critical point. In contrast, in the QSD unraveling, it undergoes an area-to-area transition with a logarithmic peak at the same critical point (so, the entanglement entropy is size-independent across the phase diagram except at the critical point). This stark difference arises because entanglement entropy is a nonlinear function of the quantum state, making it intrinsically sensitive to correlations, whereas stabilizer entropy, being expressible solely in terms of expectation values, is not.
To illustrate this point, we show one example of dynamics in Fig. 1 for a system of $N=40$ spins, in the BTC phase [$\omega_{0}=2\kappa$ in Eq. (2)]. The Lindblad dynamics is unraveled in terms of pure-state quantum trajectories, and averages are computed over $M=2000$ realizations for all considered unravelings, specified by the value of the parameter $\mu$ in Eq. (5). In panel (a), where we plot the nonstabilizerness density, the curves relative to the different unravelings are very close to each other (we have numerically verified that the difference between curves decreases by increasing $N$), whereas in panel (b) we see that the unravelings have completely different average bipartite entanglement. As expected, the magnetization along the $z$ axis, defined as $m_{z}=\langle\hat{S}_{z}\rangle/S$, is the same among all unravelings, panel (c), as for any other observable. Similar results are valid for the full distributions along quantum trajectories, as we show in Appendix A.
These findings highlight that, in a regime where quantum features are present as testified by the nonzero entanglement entropy, different physical quantities relevant for quantum information behave differently under monitored dynamics. While the stabilizer entropy remains largely unraveling-independent due to its reliance on expectation values, the entanglement entropy—being inherently sensitive to correlations—exhibits a strong dependence on the monitoring protocol. This contrast highlights the fact that quantum complexity manifests in fundamentally different ways depending on the chosen metric, as each assigns a different weight to correlations in the system.